OpenStudy (anonymous):

There's a function table and they want me to find the missing value. It looks like this: x y -2 3 -1 0 0 -1 1 2 3 I also need help figuring out what the function is between them. If you have a specific method, please share. It's really hard to grasp for me :(

OpenStudy (ghazi):

lets say your function varies like y=k*x , where k is constant now put y= 3, x=-2 to find k 3=k*-2 k=-2/3 there you have the constant now in next case you have x=1, k= -2/3 but y= ? plug in the values y=-2/3*1=-2/3

OpenStudy (ghazi):

and your required function is \[y=\frac{ -2 }{ 3 }x\]

OpenStudy (anonymous):

Thank you so much for your help!